# Maël Dumas LIFO EA 4022, Univ. Orléans, INSA Centre Val de Loire, France Anthony Perez LIFO EA 4022, Univ. Orléans, INSA Centre Val de Loire, France Mathis Rocton Algorithms and Complexity Group, TU Wien, Vienna, Austria Ioan Todinca LIFO EA 4022, Univ. Orléans, INSA Centre Val de Loire, France

### v2

@@ 3 Kernelization algorithm for Block Graph Editing
We begin this section by providing a high-level description of our kernelization algorithm. A similar technique will be applied for modification problems towards strictly chordal graphs but on the critical clique graph rather than the original one. Let us consider a positive instance $(G=(V,E),k)$ of Block Graph Editing, $F$ a suitable solution and $H=G\triangle F$. Since $|F|\leqslant k$, we know that at most $2k$ vertices of $H$ may be affected vertices. Let $A$ be the set of such vertices, $T$ the minimum induced subgraph of $H$ that spans all vertices of $A$ and $A^{\prime}$ the set of vertices of degree at least $3$ in $T$. From Lemma 5 we have $|A^{\prime}\setminus A|\leqslant 3\cdot|A|$. We will define the notion of BG-branch, corresponding to subgraphs of $G$ that induce block graphs. We will focus our interest on two types of BG-branches: the ones that are connected to the rest of the graph by only one (cut) vertex, called $1$-BG-branches, and the ones that are connected to the rest of the graph by exactly two non adjacent vertices, called $2$-BG-branches. We will prove that we can keep only the cut-vertices of $1$-BG-branches and reduce the $2$-BG-branches containing more than $k+3$ vertices. We will see that there are two kinds of connected components remaining in the graph $H\setminus(A\cup A^{\prime})$, the ones adjacent to maximal cliques of vertices of $A\cup A^{\prime}$, corresponding to critical cliques (which can be bounded to a linear number of vertices) and the ones adjacent to two non adjacent vertices of $A\cup A^{\prime}$ (which correspond to $2$-BG-branches in $H$). Since $H$ is a block graph, we know that there is at most $|A\cup A^{\prime}|=O(k)$ maximal cliques of vertices of $A\cup A^{\prime}$. From Lemma 5 we have that there are at most $4k$ connected components in $T\setminus(A\cup A^{\prime})$, and each one corresponds to a $2$-BG-branch in the graph $H$. It remains that $H\setminus(A\cup A^{\prime})$ contains a linear number of connected components, each one containing $O(k)$ vertices. Altogether, the graph $H$ contains $O(k^{2})$ vertices.
@@ 3.2 Reducing Block Graph branches
Let $G=(V,E)$ be a graph, a BG-branch is a connected induced subgraph $B$ of $G$ such that $B$ is a block graph. A vertex $v\in V(B)$ is an attachment point of $B$ if $N_{G}(v)\setminus V(B)\neq\emptyset$. A BG-branch is a $p$-BG-branch if it has exactly $p$ attachment points. We denote $B^{R}$ the subgraph of $B$ in which all attachment points have been removed. We first define a simple rule that allows to remove all the vertices of a $2$-BG-branch except for its attachment point.
@@
We now deal with the $2$-BG-branches, which require more technical skills. A $2$-BG-branch $B$ is clean if $B^{R}$ is connected and if the two attachment points are not adjacent. For a clean $2$-BG-branch $B$, a min-cut of $B$ is a set of edges $M\subseteq E(B)$ of minimum size such that the attachment points $p_{1}$ and $p_{2}$ are not in the same connected component of $B-M$ and $mc(B)$ is the size of a min-cut of $B$. A $2$-BG-branch $B$ is clean if $B^{R}$ is connected and if the two attachment points are not adjacent. For a $2$-BG-branch proper $B$, a minimum edge-cut of $B$ is a set of edges $M\subseteq E(B)$ of minimum size such that the attachment points $p_{1}$ and $p_{2}$ are not in the same connected component of $B-M$. We denote by $mc(B)$ the size of a minimum edge-cut of $B$.
@@
Assume now that $p_{1}$ and $p_{2}$ are in the same connected component in $H^{\prime}$ and let $\pi_{H^{\prime}}$ denote a shortest path between them. Note that its length is potentially $1$ if $p_{1}$ and $p_{2}$ are adjacent in $H^{\prime}$. We first consider the case where there is a path $\pi_{B}=\{p_{1}=u_{1},u_{2},\dots,u_{r}=p_{2}\}$ (not necessarily induced) from $p_{1}$ to $p_{2}$ in $B$ that still exists in $B\triangle F$. Since $B$ is a clean $2$-BG-branch of $G$ reduced by Rule 3.3, we know by the previous arguments that every cut-vertex of $B$ is in $\pi_{B}$. Note that all vertices in $\pi_{B}$ belong to the same biconnected component of $H$, which is a clique. Hence we can consider w.l.o.g. that $\pi_{B}$ is the path containing exactly $p_{1},p_{2}$ and the cut-vertices of $B$, which is the shortest path in $B$ between $p_{1}$ and $p_{2}$. Since the vertices of $\pi_{B}$ induce a path in $B$ (on at least $3$ vertices since $p_{1}$ and $p_{2}$ are not adjacent) and a clique in $B\triangle F$, then for every $i\in\{1,\dots,r-2\}$, we have $(u_{i},u_{r})\in F$. Any vertex $v\in B\setminus V(\pi_{B})$ is adjacent in $B$ to two vertices $w_{1},w_{2}\in V(\pi_{B})$ and if $v$ remains in the same biconnected component as $w_{1}$ and $w_{2}$ in $H$, then $(v,p_{1})\in F$ or $(v,p_{2})\in F$. Otherwise $F$ must contain $(v,w_{1})$ or $(v,w_{2})$. In any case, $F$ contains at least one pair for each vertex $v\in B\setminus\{u_{r-1},u_{r}\}$. It follows from the previous arguments that $|F|\geqslant|B|-2$. However, since $|B|\geqslant k+3$ we have $|F|\geqslant k+1$, contradicting that $F$ is a solution, therefore there is no such path $\pi_{B}$.
@@
Figure 4: Illustration of the case where $p_{1}$ and $p_{2}$ are in the same connected component of $H^{\prime}$ and $F$ contains a min-cut of $B$ in Lemma 21. The white circles correspond to critical cliques of $B$. The blue sets correspond to $N_{B}(P_{1})$ and $N_{B}(P_{2})$. The red dotted edges between $B_{1}$ and $B_{2}$ are edges of an edge-cut of $B$ deleted by editing.
@@
is a block graph. Let $F^{\prime}$ be the edition such that $H^{\prime}=G^{\prime}\triangle F^{\prime}$. Since $mc(B^{\prime})=mc(B)\leqslant k$, by construction $|F^{\prime}|=|F|$ , therefore $F^{\prime}$ is a $k$-edition of $G^{\prime}$. If $F$ does not contain a min-cut of $B$, then no vertices of $B^{R}$ is affected and $p_{1},p_{2}$ are not in the same connected component of $H\setminus V(B^{R})$. Let $H_{1}$ and $H_{2}$ be the connected components of $H\setminus V(B^{R})$ that contain respectively $p_{1}$ and $p_{2}$ and $H_{3}$ the remaining components of $H\setminus V(B^{R})$. By Lemma 3 the graph $H^{\prime}$ corresponding to the disjoint union of
@@ 3.3 Bounding the size of the kernel
Theorem 2.
Block Graph Editing admits a kernel with $O(k^{2})$ vertices.
@@
Let $(G=(V,E),k)$ be a reduced yes-instance of Block Graph Editing, $F$ a $k$-edition of $G$ and $H=G\triangle F$. We assume that $G$ is connected, the following arguments can be easily adapted if this is not the case by summing over all connected components of $G$. Let $A$ be the set of affected vertices of $H$. Since $|F|\leqslant k$, we have $|A|\leqslant 2k$. Let $T$ be the connected minimal induced subgraph of $H$ that spans all affected vertices of $A$, and $A^{\prime}$ the set of vertices of degree at least $3$ in $T$. By Lemma 5, $|A^{\prime}\setminus A|\leqslant 3\cdot|A|\leqslant 6k$.
@@
First, we can observe that if there is a cut-vertex $x\in V(H)$ that is not in $V(T)$, since $T$ is the connected minimal induced subgraph of $H$ that spans all affected vertices, there is a connected component $K$ of $H\setminus\{x\}$ that does not contain any affected vertex. Since $B=H[V(K)\cup\{x\}]$ is a block graph by heredity, it follows that $B$ induces a $1$-BG-branch with attachment point $x$ and can be reduced by Rule 3.3, a contradiction. It remains that every cut-vertex of $H$ lies in $V(T)$ and that every vertex of $V(H)\setminus V(T)$ is adjacent to a maximal clique of vertices or a single vertex of $V(T)$. Since the graph is reduced under Rule 3.3, the second case can not happen or else there would a $1$-BG-branch. Hence, there are two kinds of connected components in $H\setminus(A\cup A^{\prime})$, namely the ones that are adjacent to a maximal clique of vertices of $A\cup A^{\prime}$ and the ones that are adjacent to two non-adjacent vertices of $A\cup A^{\prime}$ (i.e., the ones that contain connected components of $T\setminus(A\cup A^{\prime})$). We say that these connected components are of type 1 or type 2 respectively (see Figure 5).
@@
The type 1 connected components are cliques since $H$ is a block graph and have the same neighborhood in $H$, thus they are critical cliques reduced by Rule 3.2 and contains at most $22k+$ 2 vertices. Since $T$ is a block graph, there is at most $|A\cup A^{\prime}|\leqslant 8k$ maximal cliques of vertices of $A\cup A^{\prime}$, hence there are at most $8k\cdot(2k+2)$ vertices in connected components of type 1.
@@
Now consider the connected components of type 2. We can observe that such a component $K$ together with its adjacent vertices $x,y\in A\cup A^{\prime}$ form a clean 2-$BG$-branch, and thus contains at most $32k+$ 3 vertices. By Lemma 5, $T\setminus(A\cup A^{\prime})$ contains at most $2\cdot|A|=4k$ connected components which is exactly the number of type 2 components, hence there are at most $4k\cdot(2k+1)$ vertices in connected components of type 2.
@@
We conclude that $|V(G)|=O(k^{2})$. Finally we claim that the reduced instance can be computed in polynomial time. Indeed, all reduction rules can be exhaustively applied in polynomial time (Lemmas 9 and 13).
@@
Figure 5: Illustration of a reduced instance in the proof of Theorem 2. Square nodes correspond to vertices of $T$, the ones filled in red are vertices of $A$, the ones filled in blue are vertices of $A^{\prime}\setminus A$. Circle nodes correspond to critical cliques of $G$.
@@
In this section we presented a kernel for Block Graph Editing. For the case of Block Graph Deletion, all the lemmata above are still true. In particular Lemma 11 holds by simply replacing ”$k$-edition” with ”$k$-deletion”, and the proof for the deletion variant is contained in the edition one. Using the same set of rules, we have a kernel with $O(k^{2})$ vertices for Block Graph Deletion.
Theorem 3.
Block Graph Deletion admits a kernel with $O(k^{2})$ vertices.
@@ 4 Kernelization algorithm for Strictly Chordal Editing
As in the previous section, we begin this section by providing a high-level description of our kernelization algorithm. We use the critical clique graph of strictly chordal graphs, which is a block graph (Theorem 1), to bound the number of vertices of a reduced instance. While the technique used to bound the size of the kernels for block graphs and strictly chordal graphs is similar, the structure and rules for the latter are more involved. We thus provide a stand-alone proof of this result for the sake of completeness.
Let us consider a positive instance $(G=(V,E),k)$ of Strictly Chordal Editing, $F$ a suitable solution and $H=G\triangle F$. Denote by $\mathcal{C}(H)$ the critical clique graph of $H$ as described in Definition Definition 2. Since $|F|\leqslant k$, we know that at most $2k$ critical cliques of $\mathcal{C}(H)$ may contain affected vertices. Let $A$ be the set of such critical cliques, $T$ the minimum subgraph of $\mathcal{C}(H)$ that spans all critical cliques of $A$ and $A^{\prime}$ the set of critical cliques of degree at least $3$ in $T$. Recall that $|A^{\prime}\setminus A|\leqslant 3\cdot|A|$ (Lemma 5). We will define the notion of SC-branch, corresponding to well-structured strictly chordal subgraphs of $G$. We will focus our interest on two types of SC-branches: the ones that are connected to the rest of the graph by only one critical clique, called $1$-SC-branches, and the ones that are connected to the rest of the graph by exactly two critical cliques, called $2$-SC-branches. This time the $1$-SC-branches will only be reduced to two critical cliques. The connected components of the graph $T\setminus(A\cup A^{\prime})$ correspond to subgraphs of $2$-SC-branches and the length of such $2$-SC-branches will be bounded by $k+3$. We know that there are at most $4k$ such connected components, thus there is $O(k^{2})$ critical cliques in $T$. Finally, the connected components of the graph $\mathcal{C}(H)\setminus V(T)$ correspond to $1$-SC-branches or sets of connected $1$-SC-branches. The latter case will require a specific reduction rule to be dealt with. This implies that every critical clique or maximal clique of $T$ can have a linear number of $1$-SC-branches of $\mathcal{C}(H)\setminus V(T)$ adjacent to it. Altogether, the graph $\mathcal{C}(H)$ contains $O(k^{3})$ critical cliques, and each critical clique is of size at most $k+1$, hence the graph $G$ contains $O(k^{4})$ vertices.

### v3

@@ 3 Kernelization algorithm for Block Graph Editing
We begin this section by providing a high-level description of our kernelization algorithm. A similar technique will be applied for modification problems towards strictly chordal graphs but on the critical clique graph rather than the original one. Let us consider a positive instance $(G=(V,E),k)$ of Block Graph Editing, $F$ a suitable solution and $H=G\triangle F$. Since $|F|\leqslant k$, we know that at most $2k$ vertices of $H$ may be affected vertices. Let $A$ be the set of such vertices, $T$ the minimum induced subgraph of $H$ that spans all vertices of $A$ and $A^{\prime}$ the set of vertices of degree at least $3$ in $T$. From Lemma 5 we have $|A^{\prime}\setminus A|\leqslant 3\cdot|A|$. We will define the notion of BG-branch, corresponding to subgraphs of $G$ that induce block graphs. We will focus our interest on two types of BG-branches: the ones that are connected to the rest of the graph by only one (cut) vertex, called $1$-BG-branches, and the ones that are connected to the rest of the graph by exactly two non-adjacent vertices, called $2$-BG-branches. We will prove that we can keep only the cut-vertices of $1$-BG-branches and reduce the $2$-BG-branches containing more than $k+3$ vertices. We will see that there are two kinds of connected components remaining in the graph $H\setminus(A\cup A^{\prime})$, the ones adjacent to maximal cliques of vertices of $A\cup A^{\prime}$, corresponding to critical cliques (which can be bounded to a linear number of vertices) and the ones adjacent to two non-adjacent vertices of $A\cup A^{\prime}$ (which correspond to $2$-BG-branches in $H$). Since $H$ is a block graph, we know that there is at most $|A\cup A^{\prime}|=O(k)$ maximal cliques of vertices of $A\cup A^{\prime}$. From Lemma 5 we have that there are at most $4k$ connected components in $T\setminus(A\cup A^{\prime})$, and each one corresponds to a $2$-BG-branch in the graph $H$. It remains that $H\setminus(A\cup A^{\prime})$ contains a linear number of connected components, each one containing $O(k)$ vertices. Altogether, the graph $H$ contains $O(k^{2})$ vertices.
@@ 3.2 Reducing Block Graph branches
Let $G=(V,E)$ be a graph, a BG-branch is a connected induced subgraph $B$ of $G$ such that $B$ is a block graph. A vertex $v\in V(B)$ is an attachment point of $B$ if $N_{G}(v)\setminus V(B)\neq\emptyset$. A BG-branch is a $p$-BG-branch if it has exactly $p$ attachment points. We denote $B^{R}$ the induced subgraph of $B$ in which all attachment points have been removed. We first define a simple rule that allows to remove all the vertices of a $1$-BG-branch except for its attachment point.
@@
We now deal with the $2$-BG-branches, which require more technical skills. A $2$-BG-branch $B$ is clean if $B^{R}$ is connected and if the two attachment points are not adjacent. For a clean $2$-BG-branch $B$, a min-cut of $B$ is a set of edges $M\subseteq E(B)$ of minimum size such that the attachment points $p_{1}$ and $p_{2}$ are not in the same connected component of $B-M$ and $mc(B)$ is the size of a min-cut of $B$.
@@
Assume now that $p_{1}$ and $p_{2}$ are in the same connected component in $H^{\prime}$ and let $\pi_{H^{\prime}}$ denote a shortest path between them. Note that $\pi_{H^{\prime}}$ may be of length $1$ if $p_{1}$ and $p_{2}$ are adjacent in $H^{\prime}$. We first consider the case where there is a path $\pi_{B}=\{p_{1}=u_{1},u_{2},\dots,u_{r}=p_{2}\}$ (not necessarily induced) from $p_{1}$ to $p_{2}$ in $B$ that still exists in $B\triangle F$. Since $B$ is a clean $2$-BG-branch of $G$ reduced by Rule 3.3, we know by the previous arguments that every cut-vertex of $B$ is in $\pi_{B}$. Note that all vertices in $\pi_{B}$ belong to the same biconnected component of $H$, which is a clique. Hence we can consider w.l.o.g. that $\pi_{B}$ is the path containing exactly $p_{1},p_{2}$ and the cut-vertices of $B$, which is the shortest path in $B$ between $p_{1}$ and $p_{2}$. Since the vertices of $\pi_{B}$ induce a path in $B$ (on at least $3$ vertices since $p_{1}$ and $p_{2}$ are not adjacent) and a clique in $B\triangle F$, then for every $i\in\{1,\dots,r-2\}$, we have $(u_{i},u_{r})\in F$. Any vertex $v\in B\setminus V(\pi_{B})$ is adjacent in $B$ to two vertices $w_{1},w_{2}\in V(\pi_{B})$ and if $v$ remains in the same biconnected component as $w_{1}$ and $w_{2}$ in $H$, then $(v,p_{1})\in F$ or $(v,p_{2})\in F$. Otherwise $F$ must contain $(v,w_{1})$ or $(v,w_{2})$. In any case, $F$ contains at least one pair for each vertex $v\in B\setminus\{u_{r-1},u_{r}\}$. It follows from the previous arguments that $|F|\geqslant|B|-2$. However, since $|B|\geqslant k+3$ we have $|F|\geqslant k+1$, contradicting that $F$ is a solution, therefore there is no such path $\pi_{B}$.
@@
Figure 5: Illustration of the case where $p_{1}$ and $p_{2}$ are in the same connected component of $H^{\prime}$ and $F$ contains a min-cut of $B$ in Lemma 21. The white circles correspond to critical cliques of $B$. The blue sets correspond to $N_{B}(P_{1})$ and $N_{B}(P_{2})$. The red dotted edges between $B_{1}$ and $B_{2}$ are edges of an edge-cut of $B$ deleted by the edition.
@@
is a block graph. Let $F^{\prime}$ be the edition such that $H^{\prime}=G^{\prime}\triangle F^{\prime}$. Since $mc(B^{\prime})=mc(B)\leqslant k$, by construction $|F^{\prime}|=|F|$ , therefore $F^{\prime}$ is a $k$-edition of $G^{\prime}$. If $F$ does not contain a min-cut of $B$, then no vertex of $B^{R}$ is affected and $p_{1},p_{2}$ are not in the same connected component of $H\setminus V(B^{R})$. Let $H_{1}$ and $H_{2}$ be the connected components of $H\setminus V(B^{R})$ that contain respectively $p_{1}$ and $p_{2}$ and $H_{3}$ the remaining components of $H\setminus V(B^{R})$. By Lemma 3 the graph $H^{\prime}$ corresponding to the disjoint union of
@@ 3.3 Bounding the size of the kernel
See 1
@@
Let $(G=(V,E),k)$ be a reduced yes-instance of Block Graph Editing, $F$ a $k$-edition of $G$ and $H=G\triangle F$. We assume that $G$ is connected, the following arguments can be easily adapted if this is not the case by summing over all connected components of $G$. Let $A$ be the set of affected vertices of $H$. Since $|F|\leqslant k$, we have $|A|\leqslant 2k$. Let $T_{A}$ be the connected minimal induced subgraph of $H$ that spans $A$, and $d(T_{A})$ the set of vertices of degree at least $3$ in $T_{A}$. By Lemma 5, $|d(T_{A})\setminus A|\leqslant 3\cdot|A|\leqslant 6k$.
@@
First, we can observe that if there is a cut-vertex $x\in V(H)$ that is not in $V(T_{A})$, since $T_{A}$ is the connected minimal induced subgraph of $H$ that spans all affected vertices, there is a connected component $K$ of $H\setminus\{x\}$ that does not contain any affected vertex. Since $B=H[V(K)\cup\{x\}]$ is a block graph by heredity, it follows that $B$ induces a $1$-BG-branch with attachment point $x$ and can be reduced by Rule 3.3, a contradiction. It remains that every cut-vertex of $H$ lies in $V(T_{A})$ and that every vertex of $V(H)\setminus V(T_{A})$ is adjacent to a maximal clique of vertices or a single vertex of $V(T_{A})$. Since the graph is reduced under Rule 3.3, the second case can not happen or else there would a $1$-BG-branch. Hence, there are two kinds of connected components in $H\setminus(A\cup d(T_{A}))$, namely the ones that are adjacent to a maximal clique of vertices of $A\cup d(T_{A})$ and the ones that are adjacent to two non-adjacent vertices of $A\cup d(T_{A})$ (i.e., the ones that contain connected components of $T_{A}\setminus(A\cup d(T_{A}))$). We say that these connected components are of type 1 or type 2 respectively (see Figure 6).
@@
The type 1 connected components are cliques since $H$ is a block graph and have the same neighborhood in $H$, thus they are critical cliques reduced by Rule 3.2 and contains at most $22k+$ 2 vertices. Since $T_{A}$ is a block graph, there is at most $|A\cup d(T_{A})|\leqslant 8k$ maximal cliques of vertices of $A\cup d(T_{A})$, hence there are at most $8k\cdot(2k+2)$ vertices in connected components of type 1.
@@
Now consider the connected components of type 2. We can observe that such a component $K$ together with its adjacent vertices $x,y\in A\cup d(T_{A})$ form a clean 2-$BG$-branch, and thus contains at most $32k+$ 3 vertices. By Lemma 5, $T_{A}\setminus(A\cup d(T_{A}))$ contains at most $2\cdot|A|=4k$ connected components which is exactly the number of type 2 components, hence there are at most $4k\cdot(2k+1)$ vertices in connected components of type 2.
@@
We conclude that if $(G,k)$ is a reduced yes-instance, then $|V(G)|\leqslant 24k^{2}+28k$. Therefore, given a reduced instance $(G,k)$ of Block Graph Editing, if $|V(G)|>24k^{2}+28k$ we can safely conclude that $(G,k)$ is a no-instance and return a trivial no-instance of constant size, thus proving the size of our kernel. Finally we claim that the reduced instance can be computed in polynomial time. Indeed, all reduction rules can be exhaustively applied in polynomial time (Lemmata 9 and 13).
@@
Figure 6: Illustration of a reduced instance in the proof of Theorem 1. Square nodes correspond to vertices of $T_{A}$, the ones filled in red are vertices of $A$, the ones filled in blue are vertices of $d(T_{A})\setminus A$. Circle nodes correspond to critical cliques of $G$.
@@
For the case of Block Graph Deletion, all the lemmata in this section are still true. In particular Lemma 11 holds by simply replacing ”$k$-edition” with ”$k$-deletion”, and the proof for the deletion variant is contained in the edition one. Using the same set of rules, we have a kernel with $O(k^{2})$ vertices for Block Graph Deletion. ∎
@@ 4 Kernelization algorithm for Strictly Chordal Editing
As in the previous section, we begin this section by providing a high-level description of our kernelization algorithm. We use the critical clique graph of strictly chordal graphs, which is a block graph by definition, to bound the number of vertices of a reduced instance. While the technique used to bound the size of the kernels for block graphs and strictly chordal graphs is similar, the structure and rules for the latter are more involved. We thus provide a stand-alone proof of this result for the sake of completeness.
Let us consider a positive instance $(G=(V,E),k)$ of Strictly Chordal Editing, $F$ a suitable solution and $H=G\triangle F$. Denote by $\mathcal{C}(H)$ the critical clique graph of $H$ as described in Definition 2. Since $|F|\leqslant k$, we know that at most $2k$ critical cliques of $\mathcal{C}(H)$ may contain affected vertices. Let $A$ be the set of such critical cliques, $T$ the minimum subgraph of $\mathcal{C}(H)$ that spans all critical cliques of $A$ and $A^{\prime}$ the set of critical cliques of degree at least $3$ in $T$. Recall that $|A^{\prime}\setminus A|\leqslant 3\cdot|A|$ (Lemma 5). We will define the notion of SC-branch, corresponding to well-structured strictly chordal subgraphs of $G$. We will focus our interest on two types of SC-branches: the ones that are connected to the rest of the graph by only one critical clique, called $1$-SC-branches, and the ones that are connected to the rest of the graph by exactly two critical cliques, called $2$-SC-branches. This time the $1$-SC-branches will only be reduced to two critical cliques. The connected components of the graph $T\setminus(A\cup A^{\prime})$ are paths, which correspond to subgraphs of $2$-SC-branches, and we will bound their length by $k+3$. We know that there are at most $4k$ such connected components, thus there is $O(k^{2})$ critical cliques in $T$. Finally, the connected components of the graph $\mathcal{C}(H)\setminus V(T)$ correspond to $1$-SC-branches or sets of connected $1$-SC-branches. The latter case will require a specific reduction rule to be dealt with. This implies that every critical clique or maximal clique of $T$ can have a linear number of $1$-SC-branches of $\mathcal{C}(H)\setminus V(T)$ adjacent to it. Altogether, the graph $\mathcal{C}(H)$ contains $O(k^{3})$ critical cliques, and each critical clique is of size at most $k+1$, hence the graph $G$ contains $O(k^{4})$ vertices.